OpenStudy (idku):

Guys I suck at functions, can you show me everything about them? Don't link me! Guide me, don't do the work for me! HELP!!!!

OpenStudy (anonymous):

haha, most people have question about fraction not 'function's sory

OpenStudy (idku):

Can you teach me things about functions, I know my q is broad but...

OpenStudy (anonymous):

where is your question about functions? there is alot of info...

OpenStudy (idku):

like the function translations and parabolas ellipses and etc

OpenStudy (anonymous):

f(x) means that there is an equation that has x in it and when you plug in x and solve, then the answer will be spit out

OpenStudy (idku):

spit out?

OpenStudy (anonymous):

function translations, like how does a the function of a parabola... ?

OpenStudy (idku):

Function translations (like f(x)= .... ) Parabolas Ellipses

OpenStudy (anonymous):

oh, you are talking about conic sections. and how to figure out what the graph is based on the function?

OpenStudy (idku):

Yeah, like when it gives the choices f which is the graph of the function.

OpenStudy (anonymous):

so a parabola, circle, elipse, and hyperboly are all 'conic sections' and as such, the all have the same 'standard form' but one standard from will appear different than another standard form. if the function can be written in the form 'vertex form' or 'circle form' then you can know if it is a circle or parabola or what ever.

OpenStudy (idku):

Can we start from parabolas?

OpenStudy (anonymous):

so a parabola will always have x^2, so it can be like x^2 +x +7 or it can be x^2+8 stuff like that

OpenStudy (idku):

Can you guide me through couple examples?

OpenStudy (anonymous):

mmm.. this is Big question you are asking than I had inticipated to bite off! but here is an outline of how things look: we are normally use to seeing a parabola like this: Ax^2 + Bx + C = 0, then you use the quadratic equation with the coefficeints A,B and the constant C. well, conic sections are the same idea when looking at them all together, but the coefficients are labled differently: before we just did A, B, C to lable them, now we use A with the X^2 term and D with the x term. look at these set ups I pulled off of "conics" off of "purple math" website parabola: Ax^2 + Dx + Ey = 0 circle: x^2 + y^2 + Dx + Ey + F = 0 ellipse: Ax^2 + Cy^2 + Dx + Ey + F = 0 hyperbola: Ax^2 – Cy^2 + Dx + Ey + F = 0

OpenStudy (anonymous):

here is the website http://www.purplemath.com/modules/conics.htm if you go to the part that I copied and pasted the info, there is a 'check list' to walk through to see what the function is when graphed.

OpenStudy (anonymous):

its this: " Are both variables squared? No: It's a parabola. Yes: Go to the next test.... Do the squared terms have opposite signs? Yes: It's an hyperbola. No: Go to the next test.... Are the squared terms multiplied by the same number? Yes: It's a circle. No: It's an ellipse.

OpenStudy (anonymous):

haha, thanks for the testimony purple math website is a great site though, they teach in words that can be understood. we good here? i'm gonna take my dog on a walk